(y - xy)dx + xdy = 0



1
Expert's answer
2022-01-31T09:36:25-0500

"M=(y-xy)\\
\frac{dM}{dy}=1-xy\\
N= x\\
\frac{dN}{dy}=1\\
which ~is~ not~exact\\
( \frac{dM}{dy}-\frac{dN}{dy})=f(x)\\
(1-x-1)=f(x),-x=f(x)\\
\frac{1}{N}( \frac{dM}{dy}-\frac{dN}{dy})=f(x)\\
\frac{1}{x}(-x)=f(x)\\
-1=f(x)\\
find~the~integrating~factor\\
\frac{1}{N}( \frac{dM}{dy}-\frac{dN}{dy})=f(x), I.F=e^{\int f(x)dx}\\
\frac{1}{M}( \frac{dM}{dy}-\frac{dN}{dy})=f(y), I.F=e^{\int -f(y)dx}\\
 I.F=e^{\int f(x)dx}=e^{\int -1dx}=e^{-x}\\
e^{-x}[(y-xy)dx+xdy=0]\\
(ye^{-x}-xye^{-x})dx+xe^{-x}dy=0\\
test~for~exactness\\
\frac{dM}{dy}=e^{-x}-xe^{-x},\frac{dN}{dy}=e^{-x}-xe^{-x}\\
\frac{dF}{dx}=M(x,y)\\
\int dF=\int (ye^{-x}-xye^{-x})dx\\
F=-ye^{-x}-y\int xe^{-x}dx\\
we~use~integration~by~part~to~solve~\int xe^{-x}dx\\
\int xe^{-x}dx=xe^{-x}+e^{-x}\\
F=-ye^{-x}+yxe^{-x}+ye^{-x}+Q(y)\\
F=xye^{-x}+Q(y)\\
\frac{dF}{dy}=N(x,y)\\
\int dF=\int (xe^{-x})dy\\
F=(xe^{-x})\int dy\\
F=xye^{-x}+R(x),then\\
F=xye^{-x}+Q(y)=xye^{-x}+R(x)\\Q(x)=R(x)=0\\
xye^{-x}+Q(y)=F\\
xye^{-x}=c\\
The~general~equation"